Work Progress

Mo Tu We Th Fr Sa Su
26 27  28 29 30 31  1
 2  3   4  5  6  7  8
 9 10  11 12 13 14 15
16 17  18 19 20 21 22
23 24  25 26 27 28 29
30 31   1  2  3  4  5
Mo Tu We Th Fr Sa Su
30 31   1  2  3  4  5
 6  7   8  9 10 11 12
13 14  15 16 17 18 19
20 21  22 23 24 25 26
27 28  29 30 31  1  2
Mo Tu We Th Fr Sa Su
27 28  29 30 31  1  2
 3  4   5  6  7  8  9
10 11  12 13 14 15 16
17 18  19 20 21 22 23
24 25  26 27 28 29 30
Mo Tu We Th Fr Sa Su
 1  2   3  4  5  6  7
 8  9  10 11 12 13 14
15 16  17 18 19 20 21
22 23  24 25 26 27 28
29 30  31  1  2  3  4
Mo Tu We Th Fr Sa Su
29 30  31  1  2  3  4
 5  6   7  8  9 10 11
12 13  14 15 16 17 18
19 20  21 22 23 24 25
26 27  28 29 30  1  2

← September


we leave some overnight PCRs reactions, using VF2 and VR, in order to properly analyze our clones...

at the same time, we are inoculating all those clones in order to make some minipreps.


we have done the gel of the PCR reactions, but it looks like the DNA has not amplified our samples, but the reaction has worked fine for the controls.

we repeat the PCRs and make some other controls with different DNA quantities. the gels shows some bands, but their weights do not correspond to anything we wanted... could be some other lab dwarf??

we have miniprepped the inoculations that have grown overnight. this way, we will have more DNA to sequence


we have digested the plasmid that were amplified with the PCRs with EcoRI and PstI, because we want to be sure if it is, or if it is not, what we want... at the gel we see two bands on one clone and three (??) on the other... and the weights are not the correct ones... we must have had some false positives clones...

VAdv.jpg They ARE BioBricks, as they get amplified with VR and VF2 and their miniprep yields a plasmid that corresponds to the ones we want. But, for the construction pLac-TetR-CFP, when we have digested it with EcoRI and PstI we see three bands: 2,2 kb, 2 kb and 400 bp. For the construction pTet-LacI-YFP, we see two bands: one at 3kb and the other at 2,8 kb. sadly, this does not correpond to any simple, straigh forward hipothesys we may come up with...

we do more sequencing reactions of pTet-LacI. We are 100% sure that our pLac-TetR is what is what meant to be, but with the consensus sequence of pTet-LacI still has some gaps and missed parts. We want to be sure to have the correct sequence of the comparator part.

we prepare more inoculations to purify tomorrow.


we are miniprepping and digesting the comparator device (pTet-LacI and pLac-TetR) with the fluorescences (CFP and YFP). we are purifying it and concentrating it so tomorrow we could ligate it.


this morning we have ligated pTet-LacI-YFP and pLac-TetR-CFP. we will transform it, cross fingers (again) and hope that tomorrow we have transformants


this morning we have had eight transformants, four of each clone. we have inoculated them and we will miniprep it...


this morning we have miniprepped yesterday's inoculations. we are screening the clones in order to see if any of them is a successful ligation. we are doings some PCRs with the vectors VR and VF2 and we have digested the plasmids with EcoRI. at the gels we're looking for the correct size bands.

the gels do not give us much information... the PCR has not amplified (we have a band corresponding to the BioBricks only at the pTet-LacI control) and the digestion gel has too much DNA quantity and is not clear enough. it looks like we DO have pTet-LacI-YFP, and we DO NOT have pLac-TetR-CFP...


we have miniprepped the inoculations we prepared yesterday...

we have repeated the digestion gel and have seen a single band at the pTet-LacI-YFP and two bands at the pLac-TetR-CFP lanes... we decide to take as valid pTet-LacI-YFP from yerterday and pLac-TetR-CFP from 3 of October.

today we have repeated the PCRs changing the DNA quantities, the Tm and the time the polymerase is elongating. We want to amplify the BioBricks of two selected clones so to be sure that they are what we want.


the gels from yesterday have yield no result... looks like the problem may not be the Tm or the time of elongation, but the vial of polymerase used...

we prepare more overnight PCR reactions in order to solve this problem...


today we have received (finally!) the sequence of our last sequenciation... we are going to work on them looking for a good consensus sequence. We are 100% sure that we have pTetR-LacI as it was meant to be.

on the PCRs, we have seen that we only have amplification when we use 300 ng of template DNA. Looks like this is not normal, but the primers work as well as the Tm at 55ÂșC and the elongation time of 3'30...


VAdv.jpg we are going to make a slight change on the experimental design. we have had many, many problems ligating the comparator parts with the fluorescence. we are miniprepping pTet-Lac and pLac-Tet and we plan to digest it as an insert to assemble it backwards of pOmpR and, afterwards, try to assemble the fluorescences at the back.


we have finally set up the protocol for the PCR of our big constructs. and we are doing some PCRs to finally see if we DO have the good sizes for pLac-Tet-CFP and pTet-Lac-YFP (but, the problem is that we have a 400 pb band when we digest it!)


we are ligating pTet-Lac and pLac-Tet with the promoters pOmpR, in order to make the Promoter calibrator device.

at the same time, we are going to take some constructs that seem to be pLac-Tet-CFP and pTet-Lac-YFP with the same pOmpR.


we have had many colonies for the constructs pOmpR-pLac-Tet and pOmpR-Tet-Lac, as well as pOmpR-pLac-Tet-CFP and pOmpR-pTet-Lac-YFP. we have inoculated them and will run some gels in order to indentify which clones are the good ones.

this afternoon we will inoculate the sequenced parts we have, the ones that build up the comparator. so to send them to the Registry.


we have found a bunch of clones that correspond to pOmpR-pLac-Tet. we are going to digest them and assemble the fluorescences at its back.

VAdv.jpg the constructions we already have with the fluorescence proteins, once they are cut, have a weird running behaviour on a gel... we hope that the ligations we'll do today will bring clones that have a correct digestion pattern...


we are digesting the next ligation's parts (pOmpR-pLac-Tet as a vector and CFP as an insert), and purifying them...

with pOmpR-pTet-Lac we are finding that all the colonies we have are false positives clones: vectors that have ligated themselves without the BioBricks we want to assemble.

VAdv.jpg by now, Arnau is the only one to believe that we might have pOmpR-pLac-Tet-CFP and pOmpR-pTet-Lac-YFP from the old ligations. he wants to run the bacteria through a flow citometer, in order to be sure that they are not what we want.


today we have ligated pOmpR-pLac-Tet with CFP and we have transformed it.

as we want to have pOmpR-pTet-Lac, we are digesting and purifying each part.


Raul has come this morning to take the Petri dishes out of the oven... surprisingly, he has found many, many colonies (hundreds!)...


This afternoon we have inoculated twenty clones ao to screen tomorrow if we have any good construct...


We have miniprepped the clones and ran a gel... only to see that the plasmids that supposedly bear pOmpR-pLac-Tet-CFP are of the same size than the control pOmpR-pLac-Tet.

we want to have pOmpR-pTet-Lac, so we have ligated pOmpR as a vector and pTet-Lac as an insert... let's hope that tomorrow we have some colonies...

this afternoon we have had the first drill of the presentation we will give at Boston.


there are some colonies of pOmpR-pTet-Lac. We are going to inoculate it, miniprep it and run a gel in order to screen for a good clone.

Bye, bye BioBrick

we have ioculated the things we want to send to the registry...


we have miniprepped the possible clones pOmpR-pTet-Lac and ran a gel... only to see that we have not a single good colony... looks like our lab dwarfs do not like last-minute works... either last ligations of pOmpR-pLac-Tet-CFP or pOmpR-pTet-Lac have worked...

we have miniprepped the parts that will be sent to Boston...


we have sent our BioBricks to the MIT... let's hope they arrive in good conditions to Meagan!

we are going to meet today so to finish the presentation and the poster for Boston.


we are finishing the wiki, the poster and the presentation.

we are tracking the Biobricks: this morning they were already in Philadelphia!